Consolidated Edison has just paid an annual dividend of $3 per share. If the expected growth rate for Con Ed is 10%, and your required rate of return is 16%, how much are you willing to pay for this stock?

$55

$50

$46.50

none of the above
Answer #1

This question requires application of constant growth dividend discount model according to which: Po - Divi T-9 Po = Price of

Div1 = $3 * (1 + g) = $3 * (1 + 10%) = $3.3

3,3 ΓOO.16 – 0.10

Po = 55 - --> Answer